ganeshie8 (ganeshie8):

show that \[\large x^ty^{1-t}\le tx+(1-t)y\] \(x,y\gt 0\) and \(0\lt t\lt 1\)

Parth (parthkohli):

I can't help but notice the similarity between this and the parametric form of a complex number.

ganeshie8 (ganeshie8):

what parametric form.. polar ?

Parth (parthkohli):

Nah. Never mind.

ganeshie8 (ganeshie8):

I think you're referring to equation of straight line between two complex numbers in parametric form.. then there is some similarity yeah :)

OpenStudy (anonymous):

Take logarithm ofboth sides, note that the log function is concave and rest is simply the statement of Jensen's inequality.

OpenStudy (anonymous):

That is, \(t\log x + (1-t)\log y \leq \log (tx + (1-t)y)\)

ganeshie8 (ganeshie8):

Brilliant!

OpenStudy (jtvatsim):

I think I found an "intuitive" proof using some basic Calculus... will post a pdf soon. :)

ganeshie8 (ganeshie8):

Please.. :) that geometric proof using Jensen's inequality is pretty neat, but im not really sure how popular that inequality is...

ganeshie8 (ganeshie8):